IN 2012, AZANIA HAD A NOMINAL GDP OF $6.6 MILLION. INFLATION RATE WITH 2011 AS THE BASE YEAR WAS 105. IN 2013 THE COUNTRY'S NOMINAL GDP INCREASED TO $6.9 MILLION WITH INFLATION RATE OF 120 ON THE BASIS OF THE SAME 2011 BASE YEAR. CALCULATE THE GROWTH OF REAL GDP BETWEEN 2012 AND 2013.
1
Expert's answer
2014-11-25T09:42:23-0500
GDPn(2012) = $6.6m, CPI(2012) = 105, GDPn(2013) = $6.9m, CPI(2013) = 120
Real GDP = Nominal GDP/CPI*100
Real GDP(2012) = 6.6/105*100 = $6.29 million
Real GDP(2012) = 6.9/120*100 = $5.75 million
Real GDP growth rate = (5.75 - 6.29)/6.29*100% = -8.6%